# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ III.3 Non-zero $J_{\text{sd}}$ coupling and AFM order
We now turn to the behavior close to the interface. The staggered densities arise due to magnitude differences between the eigenfunctions and, therefore, also couple to the Friedel oscillations of the individual eigenfunctions. In the non-zero $J_{\text{sd}}$ case, this constitutes a coupling between Friedel oscillations and the staggered magnetization, causing an interface-induced modulation of the bulk staggered order. In Fig. 6(a), we plot the spin-up and spin-down densities. In Fig. 6(b), we plot the staggered and unit cell magnetizations, where the latter is averaged over all unit cell sites. While the unit cell magnetization approximately vanishes in the center of the ribbon, significant deviations arise in the vicinity of the interfaces originating from the staggered magnetization. Since the effective continuum model of AFM considered in Sec. II.1 does not capture the uncompensated nature of the interface, it omits this type of coupling between Friedel oscillations and magnetization.
Figure 7: (a) The spin-up and spin-down site-resolved densities for the system in Fig. 6 are Fourier transformed to obtain the frequencies of the density oscillations. The dominant frequency is denoted $2k_{F,1}$. By considering the spectral function integrated over all y-momenta and evaluated at the Fermi level (c), the dominant frequency can be seen to correspond closely to the peak in the distribution of $k_{F}^{x}$, i.e. x-components of the Fermi vector of states at the Fermi level. The additional peak in the spectrum in (a) arise because the magnitude of the $\pi$ staggered order itself is modulated by the density oscillations, giving rise to a $|2k_{F}\pm\pi|$ peak. (b) The site-resolved magnetization contains the same frequencies as those of the underlying spin-densities while the dominant frequency of the unit cell magnetization (d) is $2k_{F,1}^{*}$, corresponding to the reduced Brillouin zone representation of $2k_{F,1}$, caused by aliasing.
@@
We point out that the Fermi momentum itself is spin-independent in the AFM and one thus do not expect any magnetization contribution from Fermi surface mismatch, which the continuum model confirms. The frequency of the density and magnetization oscillations are shown in detail in Fig. 7. Notice in particular that we observe the unit cell magnetization to oscillate at $2k_{F}$ exactly due to the effective coupling between the $\pi$-order and Friedel oscillations. Therefore, one of the ways to achieve a surface-induced magnetization in continuum models would be to introduce spin-dependent boundary conditions, which would correspond to a different environment of spin-up and spin-down sites in the lattice model.
@@
Following this reasoning, the emergence of uncompensated edge magnetization in itinerant electrons in an AFM depend critically on the unit cell geometry and the nature of the interface. The interface-induced magnetization arise only if the up/down spins interact differently with the interface, creating a spin-dependent environment which does not vanish upon averaging over the interface. To achieve this, an uncompensated interface is typically needed. For a compensated interface, there will still be relative shifts between the unit cell eigenfunctions, but the shifts will vanish when summing over a unit cell as both spins interact with the interface symmetrically. Along these lines, we also expect this effect to be vulnerable towards surface structure and roughness. As the magnetization is induced solely through the one lattice constant shift, a roughening of the surface could cancel out the spin-dependent interface environment.
@@ III.4 Phenomenological model for magnetization oscillations in antiferromagnets
Figure 8: Spin-up and spin-down densities (a) for an AM with $J_{\text{sd}}=t$, $\varepsilon_{2(3)}=\pm\delta$ and $\mu=-3.5t$. Due to the split non-magnetic potentials, the ribbon is no longer symmetric with respect to the combined action of time-inversion and spatial inversion. The interface-induced magnetization shown in (b) is therefore not symmetric around the ribbon center, but depict unique characteristics on the two interfaces. In contrast with the AFM case, a beating pattern arise in the unit cell magnetization, pointing towards several different frequencies being present. Note that we have plotted the unit cell magnetization magnitude multiplied by a factor 5 for clarity.
Figure 9: (a) The spin-up and spin-down site-resolved densities for the system in Fig. 8 are Fourier transformed to obtain the frequencies of the density oscillations. In contrast with the AFM (Fig. 7), the spin-up and spin-down densities depict different dominant frequencies. Note the presence of the additional peaks due to the magnitude modulation of the staggered order. The spin-asymmetry in the density frequencies can be attributed to the distribution of the spin-dependent $k_{F,\sigma}^{x}$ in the eigenfunctions at the Fermi level, obtained by a partial integration of the bulk spectral function (c). Notice also the inset in (c) showing the difference between the spin-up and spin-down spectral function at the Fermi level. The frequencies present in the site-resolved magnetization are shown in (b) and a general observation is that the spin-dependence of the density oscillations give rise to a magnetization which oscillates with all underlying frequencies. Finally, we plot the Fourier transform of the unit cell magnetization in (d) and we observe two distinct peaks, ${k}_{F}^{*\uparrow}$, $k_{F}^{*\downarrow}$, corresponding to the dominant $k_{F}^{\uparrow}$, $k_{F}^{\downarrow}$ from (c) in the reduced Brillouin zone representation due to aliasing.
@@ III.5 Non-zero $J_{\text{sd}}$ coupling and AM order
The advantage of the microscopic model is that it allows us to continuously introduce the altermagnetic (AM) order on top of the underlying antiferromagnetic (AFM) lattice. By keeping the s-d coupling and reintroducing the split potentials at the nonmagnetic sites $\varepsilon_{2(3)}=\pm\delta$, $\mathcal{PT}$ symmetry is broken and the AFM becomes an AM with spin-split bands.
@@
The nature of the magnetization in the altermagnet is quite similar to the antiferromagnet discussed in Sec. III.3. The s-d coupling gives rise to $\pi$-periodic density waves, the ones for spin-up and spin-down species being out of phase, see Fig. 8(b). The presence of the $\varepsilon_{2(3)}$ potentials only complicates the picture somewhat by lifting the equivalence between density oscillations on the two rows. As such, the interface-induced magnetization discussed in Sec. III.3 carries over to the AM as well. The uncompensated interface introduces an effective coupling between the staggered magnetization and Friedel oscillations, causing uncompensated unit cell magnetization to emerge in the vicinity of the interface.
@@
However, a key distinguishing factor between the interface-induced magnetization in AFMs and AMs is the spin-dependence of the Fermi surface in the latter. The Friedel oscillations in AMs are themselves spin-dependent as the distribution of $k_{F}^{x}$ momenta on the Fermi surface is different for the two spin-species, see Fig. 9. This gives rise to an additional source of magnetization as the spin-up and spin-down densities oscillate with different frequencies, giving rise to an additional magnetization contribution. This is the effect found within the effective continuum model in Sec. II. As these oscillations originate with the Fermi surface structure, we expect it to be more robust towards surface roughness and structure. Any source of Friedel oscillations in a system, be it an interface or an impurity  [^3] Both spinless and spinfull impurities are expected to lead to the magnetization oscillations in altermagnets., is thus expected to be sources of uncompensated magnetism in altermagnets. The spin-dependent Friedel oscillations have also recently been shown to influence the interaction between impurities that are magnetic [57, 58].
@@ IV Summary
In the lattice model, considered in Sec. III, this picture is different. Being able to describe microscopic magnetic moments, the lattice model provides access to more details. We found that there is another contribution to interface-induced magnetization originating from an effective coupling between Friedel oscillations and staggered magnetic order; this contribution is qualitatively similar for antiferromagnets, see Figs. 6 and 7, and altermagnets, see Figs. 8 and 9. This coupling is the only source of edge magnetization for the antiferromagnet within our lattice model, and we argue that it corresponds to an effective spin-dependent boundary condition in the analytical continuum model. The corresponding phenomenological model is discussed in Sec. III.4.
The coupling between Friedel oscillations and the staggered order depends critically on the lattice geometry and yields a zero net interface magnetization for compensated AFM and AM interfaces. As such, this contribution is expected to be sensitive to surface defects and roughness. The Fermi-surface contribution inherent to altermagnets requires only the spin-split structure of the Fermi surface and does not rely on the microscopic structure. Any non-magnetic sources of density oscillations, caused by, e.g., interfaces or impurities, are therefore expected to give rise to local uncompensated magnetization, which is an important aspect to be aware of for the implementation altermagnet-based spintronics devices. The observation of robust magnetization induced by non-magnetic defects could also provide an alternative way to probe altermagnets.

### v2

@@ III.3 Non-zero $J_{\text{sd}}$ coupling and AFM order
Figure 7: Spin-up and spin-down densities are shown in (a) for an AFM with $J_{\text{sd}}=t$, $\mu=-2.75t$. The spin-dependent $J_{\text{sd}}$ gives rise to highly spin-dependent occupations on the magnetic sites $\alpha=1,4$. (b) The magnitude differences in the underlying eigenfunctions gives rise to a staggered magnetization (green and orange). The magnetization is out of phase between the two rows causing the total unit cell magnetization (black) to approach zero in the ribbon center. Despite no spin-dependence in the Friedel oscillations, we still observe uncompensated magnetization (black) in the vicinity of the interface. This is caused by an effective coupling between the staggered order and the spin-independent Friedel oscillations, not captured in the continuum model. Note that we here plot the magnitude of the unit cell magnetization multiplied by a factor 5 for clarity.
@@
We now turn to the behavior close to the interface. The staggered magnetization arises due to magnitude differences between the spin-up and down densities. As the density waves for spin-up and down electrons are shifted relative to the interface, an effective coupling between Friedel oscillations and the staggered magnetization emerges, causing an interface-induced modulation of the bulk staggered order.
@@
In Fig. 7(a), we plot the spin-up and spin-down densities. In Fig. 7(b), we plot the staggered (row-resolved) and unit cell magnetizations, where the latter is averaged over all unit cell sites. While the unit cell magnetization practically vanishes in the center of the ribbon, significant deviations from the compensated bulk order arise in the vicinity of the interfaces. As the effective continuum model of AFM considered in Sec. II.1 does not capture the uncompensated nature of the interface nor the staggered order, it omits this type of coupling between Friedel oscillations and magnetization.
@@
We point out that the Fermi momentum itself is spin-independent in the AFM and one thus do not expect any magnetization contribution from Fermi surface mismatch, which the continuum model confirms. The frequency of the density and magnetization oscillations are shown in detail in Fig. 9. Notice in particular that we observe the unit cell magnetization to oscillate at $2k_{F}$ [9(d)] exactly due to the effective coupling between the $\pi$-order and Friedel oscillations. Therefore, one of the ways to achieve a surface-induced magnetization in continuum models would be to introduce spin-dependent boundary conditions, which would correspond to a different environment of spin-up and spin-down sites in the lattice model.
@@
Following this reasoning, the emergence of uncompensated edge magnetization for itinerant electrons in an AFM depend critically on the unit cell geometry and the nature of the interface. The interface-induced magnetization arise only if the up/down spins interact differently with the interface, creating a spin-dependent environment which does not vanish upon averaging over the interface. To achieve this, an uncompensated interface is typically needed. For a compensated interface, there will still be relative shifts between the spin-up and down density waves, but the shifts will vanish when summing over a unit cell as both spins interact with the interface symmetrically. Along these lines, we also expect this effect to be vulnerable towards surface structure and roughness. As the magnetization is induced solely through the one lattice constant shift, a roughening of the surface could cancel out the spin-dependent interface environment. This is discussed in detail in Sec. III.6.
@@ III.5 Non-zero $J_{\text{sd}}$ coupling and AM order
Figure 8: Spin-up and spin-down densities (a) for an AM with $J_{\text{sd}}=t$, $\varepsilon_{2}=0$, $\varepsilon_{3}=100t$ and $\mu=-2.75t$. Due to the split non-magnetic potentials, the ribbon is no longer symmetric with respect to the combined action of time-inversion and spatial inversion. The interface-induced magnetization shown in (b) is therefore not symmetric around the ribbon center, but depict unique characteristics on the two interfaces. Note that we have plotted the unit cell magnetization magnitude multiplied by a factor 5 for clarity.
@@
The advantage of the microscopic model is that it allows us to introduce the altermagnetic (AM) order on top of the underlying antiferromagnetic (AFM) lattice. By keeping the s-d coupling and setting the on-site potentials $\varepsilon_{2}=0$, $\varepsilon_{3}=100t$, $\mathcal{PT}$ symmetry is broken and the AFM becomes an AM with spin-split bands (see the bandstructure in Fig. 4).
@@
The nature of the magnetization in the altermagnet is quite similar to the antiferromagnet discussed in Sec. III.3. The s-d coupling gives rise to $\pi$-periodic density waves, the ones for spin-up and spin-down species being out of phase, see Fig. 8(b). The presence of the $\varepsilon_{3}$ potential only complicates the picture somewhat by lifting the equivalence between density oscillations on the two rows. As such, the interface-induced magnetization discussed in Sec. III.3 carries over to the AM as well. The uncompensated interface introduces an effective coupling between the staggered magnetization and Friedel oscillations, causing uncompensated unit cell magnetization to emerge in the vicinity of the interface.
@@
However, a key distinguishing factor between the interface-induced magnetization in AFMs and AMs is the spin-dependence of the Fermi surface in the latter. The Friedel oscillations in AMs are themselves spin-dependent as the distribution of $k_{F}^{x}$ momenta on the Fermi surface is different for the two spin-species, see Fig. 10(a) and (c). This gives rise to an additional source of magnetization as the spin-up and spin-down densities oscillate with different frequencies which also leaves an impact on the unit cell magnetization in Fig. 10(d). This is the effect found within the effective continuum model in Sec. II. As these oscillations originate with the Fermi surface structure, we expect it to be more robust towards surface roughness and structure. Any source of Friedel oscillations in a system, be it an interface or an impurity  [^3] Both spinless and spinfull impurities are expected to lead to the magnetization oscillations in altermagnets., is thus expected to be sources of uncompensated magnetism in altermagnets. The spin-dependent Friedel oscillations have also recently been shown to influence the interaction between impurities that are magnetic [60, 61].
We point out that unit cell geometry and the AM Fermi surface is intrinsically linked. In order to observe spin-dependent Friedel oscillations, the AM Fermi surface should be oriented in a way such that the distribution of Fermi vector components normal to the interface differs between the two spin species. Due to the link with lattice geometry, however, such an orientation would typically correspond to a unit cell geometry which has an inherent asymmetry between spin-up and down species leading to an uncompensated interface. As such, it is difficult to deconvolute the effects of the particular interface configuration which is highly important for the AFM order, from that of the Fermi surface asymmetry expected to be important in AM. To disentangle magnetism arising from the interface from that arising from bulk properties, we turn to the role of the interface polarization and surface defects.
III.6 The role of interface roughness
Figure 9: (a) The spin-up and spin-down row-resolved densities for the AFM in Fig. 7 are Fourier transformed to obtain the frequencies of the density oscillations. The dominant frequency is denoted $2k_{F,1}$. By considering the spectral function integrated over all y-momenta and evaluated at the Fermi level (c), the dominant frequency can be seen to correspond closely to the peak in the distribution of $k_{F}^{x}$, i.e. x-components of the Fermi vector of states at the Fermi level. The additional peak in the spectrum in (a) arise because the magnitude of the $\pi$ staggered order itself is modulated by the density oscillations, giving rise to a $|2k_{F}\pm\pi|$ peak. (b) The row-resolved magnetization contains the same frequencies as those of the underlying spin-densities while the dominant frequency of the unit cell magnetization (d) is $2k_{F,1}^{*}$, corresponding to the reduced Brillouin zone representation of $2k_{F,1}$, caused by aliasing.
Figure 10: (a) The spin-up and spin-down row-resolved densities for the AM in Fig. 8 are Fourier transformed to obtain the frequencies of the density oscillations. In contrast with the AFM (Fig. 9), the spin-up and spin-down densities depict different dominant frequencies. Note the presence of the additional peaks due to the magnitude modulation of the staggered order. The spin-asymmetry in the density frequencies can be attributed to the distribution of the spin-dependent $k_{F,\sigma}^{x}$ in the eigenfunctions at the Fermi level, obtained by a partial integration of the bulk spectral function (c). Notice also the inset in (c) showing the difference between the bulk spin-up and spin-down spectral function at the Fermi level. The frequencies present in the row-resolved magnetization are shown in (b) and a general observation is that the spin-dependence of the density oscillations give rise to a magnetization which oscillates with all underlying frequencies. Finally, we plot the Fourier transform of the unit cell magnetization in (d) and we observe two distinct peaks, ${k}_{F}^{*\uparrow}$, $k_{F}^{*\downarrow}$, corresponding to the dominant $k_{F}^{\uparrow}$, $k_{F}^{\downarrow}$ from (c) in the reduced Brillouin zone representation due to aliasing.
Figure 11: The unit cell magnetization is shown for an AFM (a) and an AM (b) as function of defect density in the first and last unit cell on the ribbon. The system parameters are $N_{x}=76$, $N_{y}=100$, $J_{\text{sd}}=t$, $\varepsilon_{2}=0$, and $\varepsilon_{3}=100t$ (the two latter only relevant for the AM). The filling level is $\mu=-2.75t$. For each defect concentration, the specific percentage of sites in the first and last unit cell is randomly made inaccessible by adding an on-site potential of $100t$. For each particular defect density, the unit cell magnetization is averaged over 5 randomized defect configurations. The ribbon magnetization is observed to be affected by the density of defects and for the AFM, the oscillations of magnetization almost vanish when $30\%$ of interface cell sites are a vacancy. While defects have an effect on the AM as well, the suppression of the magnitude of oscillations is not as drastic as in the AFM.
Figure 12: The relation between the regular and inverted ribbon is shown in (a). We plot the sum $\langle m_{z}\rangle+\langle m_{z}^{\text{inv}}\rangle$ (b) and (d), and difference $\langle m_{z}\rangle-\langle m_{z}^{\text{inv}}\rangle$ (c) and (e) for the same AFM (b) and (c) and AM (d) and (e) as in Fig. 11. For the AFM, the sum $\langle m_{z}\rangle+\langle m_{z}^{\text{inv}}\rangle$ in (b) is zero, independent of defect density, indicating the surface origin of the magnetization. In the AM, the sum $\langle m_{z}\rangle+\langle m_{z}^{\text{inv}}\rangle$in (d) is non-zero and quite robust against surface defects, indicating that the magnetization is not purely initiated by the particular configuration of localized moments at the interface.
Our analysis so far has assumed atomically flat interfaces, but in real samples, the interfaces are commonly rough. The effect of such roughness can be investigated in our lattice model by making the interface jagged, which is achieved by adding a strong onsite potential on random lattice sites neighboring the vacuum interface. The interface roughness analysis allows us to make stronger contact with experimental measurements, since we are able to predict if the interface-induced magnetization is robust toward such disorder in the antiferromagnetic and altermagnetic case.
In Fig. 11, the unit cell magnetization is shown for an AFM (a) and AM (b) when interface roughness is included. The presence of defects break translational invariance in the y-direction. We account for this by considering a ribbon of both finite width and length, with periodic boundary conditions between the first and last lattice site along the interface. The rough interface is modelled by inserting vacancies, described by an $100t$ on-site potential, at random locations in the first and last layers of unit cells on the ribbon. We plot the magnetization averaged over all rows in the periodic direction as a function of the density of defects and observe that the magnetization is indeed affected by the roughness, both for the AM and AFM. For example, the magnetization oscillations in the AFM are almost completely suppressed for a defect density of approx. 30 Ṫhis is understood as destructive interference caused by the oscillating magnetization acquiring a row-dependent phase shift due to the local defect configuration. In the AM, the magnetization shown in the lower panel of Fig. 11 is not suppressed to the same extent. Based on our previous analysis, we would indeed expect the magnetization in the AFM to be more sensitive towards surface disorder, but it is difficult to draw a firm conclusion based on the magnetization profiles in the AFM and AM by themselves.
Therefore, in order to demonstrate the difference between the AFM and AM response to surface disorder, we introduce the concept of an inverted ribbon which has the same Fermi surface but different lattice structure (see Fig. 12(a)). Consider the ribbon as it is presented in Fig. 1. If one reverses the direction of the localized moments as well as switch $\varepsilon_{2}$ and $\varepsilon_{3}$ (the latter is only relevant for AM), a related ribbon can be obtained with the same “bulk” environment (the same Fermi surface), but with different boundary conditions. In particular, the polarization of the boundaries has been reversed. In the following, we denote the magnetization of this inverted ribbon as $\langle m_{z}^{\text{inv}}\rangle$. The purpose of introducing the inverted ribbon is to separate the magnetization contribution arising from the interface and that arising due to intrinsic spin-asymmetries in the Fermi surface. The former should be highly surface dependent, while the latter is expected to be more robust toward interface roughness.
In Fig. 12, we plot the sum [(b) and (d)] and difference [(c) and (e)] between the magnetization of the original ribbon, $\langle m_{z}\rangle$, and the inverted ribbon $\langle m_{z}^{\text{inv}}\rangle$. By taking the sum, one should be left with the magnetization contribution which does not show a perfect inversion under a reversal of the interface polarization. Along the same lines, the difference should give exactly the surface-dependent contribution. In the case of an AFM, the sum is zero, see Fig. 12(b), independent on surface defects. Based on our previous analysis, the origin of the interface-induced magnetization in the AFM is solely due to the spin-asymmetric environment close to the interface. Upon a reversal of the interface polarization, it is reasonable that the ribbon magnetization inverts as well, giving rise to a perfect cancellation. This is exactly what we observe in Fig. 12(b). For the AFM, the difference $\langle m_{z}\rangle-\langle m_{z}^{\text{inv}}\rangle$ is highly sensitive to the surface disorder, see Fig. 12(c), supporting our statement that the magnetization in AFM is less robust toward surface roughness.
For the AM, however, the magnetizations of the ribbon and its inverted counterpart do not cancel out, but leave a non-zero remainder in the sum in Fig. 12(d). As has been discussed previously, the altermagnet should have an additional contribution to the interface magnetization arising from the spin-dependence of the Friedel oscillations. This asymmetry arises from the Fermi surface and should not be dependent on the exact realization of boundary conditions or surface defects. Indeed, as one can see in Figs. 12(d) and 12(e), the sum $\langle m_{z}\rangle+\langle m_{z}^{\text{inv}}\rangle$ is less affected by the surface disorder than the difference $\langle m_{z}\rangle-\langle m_{z}^{\text{inv}}\rangle$. We argue that this analysis indicates that the interface-induced magnetization in altermagnets is more robust towards surface-defects and the particular configuration of localized moments at the interface.
We note that the role of defects in an altermagnetic candidate MnTe was recently highlighted, albeit in a different context, in Ref. [62].
@@ IV Summary
In the lattice model, considered in Sec. III, this picture is different. Being able to describe microscopic magnetic moments, the lattice model provides access to more details. We found that there is another contribution to interface-induced magnetization originating from an effective coupling between Friedel oscillations and staggered magnetic order; this contribution is qualitatively similar for antiferromagnets, see Figs. 7 and 9, and altermagnets, see Figs. 8 and 10. This coupling is the only source of edge magnetization for the antiferromagnet within our lattice model, and we argue that it corresponds to an effective spin-dependent boundary condition in the analytical continuum model. The corresponding phenomenological model is discussed in Sec. III.4.
The coupling between Friedel oscillations and the staggered order depends critically on the lattice geometry and yields a zero net interface magnetization for compensated AFM and AM interfaces. As we show in Sec. III.6, this contribution is expected to be sensitive to surface defects and roughness. The Fermi-surface contribution inherent to altermagnets requires only the spin-split structure of the Fermi surface and does not rely on the microscopic structure of the interface. Any non-magnetic sources of density oscillations, caused by, e.g., interfaces or impurities, are therefore expected to give rise to local uncompensated magnetization, which is an important aspect to be aware of for the implementation altermagnet-based spintronics devices. The observation of robust magnetization induced by non-magnetic defects could also provide an alternative way to probe altermagnets.